OpenStudy (anonymous):

What are the square roots of 64 over 144 ?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

A.-32/72 and 32/72 B.-8/12 and 8/12 C.-4/9 and 4/9 D.-4/12 and 4/12

OpenStudy (anonymous):

By square roots, do you mean 64^(1/2), 64^(1/3)... Because 64^(1/2) = 8, and 8/144 = 1/18

OpenStudy (anonymous):

I'm still not understanding

OpenStudy (kendricklamar2014):

\[\sqrt{64} = 8\]

OpenStudy (kendricklamar2014):

\[\sqrt{144} = 12\]

OpenStudy (kendricklamar2014):

What do you think the answer is @TreyYoung

OpenStudy (anonymous):

D?

OpenStudy (kendricklamar2014):

\[No, ~Try~ Again\]

OpenStudy (kendricklamar2014):

Remember: \[8/12\]

OpenStudy (kendricklamar2014):

@TreyYoung

OpenStudy (anonymous):

c

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (kendricklamar2014):

\[Which ~answer ~has ~8/12\] @TreyYoung